OpenStudy (anonymous):

How much would $200 invested at 6% interest compounded annually be worth after 6 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

283.70

OpenStudy (anonymous):

The formula is A=P(1+r/n)^(nt) where A is the final amount, P=principal or starting amount, r= rate as a decimal. n= number of compounding periods per year t=number of years

OpenStudy (anonymous):

i know he equation but each time i do it my answer is always wrong

OpenStudy (anonymous):

meant *the

OpenStudy (anonymous):

r = 0.06 n = 1 t = 6 P = 200 A = 200(1 + 0.06)^6 = 200(1.06)^6

OpenStudy (anonymous):

A=200(1+0.06)^6 A=200(1.06)^6 A=200(1.4185)=283.70

OpenStudy (anonymous):

Make sure you follow the order of operations

OpenStudy (anonymous):

How much would $500 invested at 6% interest compounded monthly be worth after 4 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

500(1 + 0.06/12)^(12*4) = 500(1.005)^48

OpenStudy (anonymous):

jamesm, your a genius

OpenStudy (anonymous):

How much would $300 invested at 4% interest compounded monthly be worth after 8 years? Round your answer to the nearest cent. Do not include units in your answer.

OpenStudy (anonymous):

300*(1 + 0.04/12)^(12*8) lol

OpenStudy (anonymous):

i am not a genius, i just know the formula

OpenStudy (anonymous):

412.92

OpenStudy (anonymous):

How much would $500 invested at 5% interest compounded annually be worth after 10 years? Round your answer to the nearest cent. Do not include units in your answer. sorry to bother yall but this is the last one

OpenStudy (anonymous):

skittlzs, not to be rude, but, have you learned how to do these problems by now?

OpenStudy (anonymous):

no, every time i i try it then comare to yall my answers r somehow way off i do the steps but i dont know wat im doing wrong

OpenStudy (anonymous):

814.45

OpenStudy (anonymous):

well, have a go at the latest one. A = P(1 + r/n)^(nt)

OpenStudy (anonymous):

i got 819.30

OpenStudy (anonymous):

how? tell me what you put into the calculator

OpenStudy (anonymous):

wat am i doing wrong?

OpenStudy (anonymous):

i do the parenthesis first, giving me 1.025. then i multiple 500*(1.025)^20. giving me 819.30

OpenStudy (anonymous):

that's because it's not 1.025. it's 1.05 and it's not 20. it's 10.

OpenStudy (anonymous):

ehhhhhhh? how?!?!

OpenStudy (anonymous):

because it's compounded annually, so n = 1 t = 10 r = 0.05, P = 500

OpenStudy (anonymous):

well i feel stupid. i put n=2

OpenStudy (anonymous):

jamesm, i appreciate your help. i really do. hope to see you around and maybe one day i can repay the favor